OpenStudy (anonymous):

a central pivot irrigation system is watering a circular field with radius 120 feet. The system rotates pi/3 radians in one hour. what area(to the nearest square foot) is watered in one hour?

OpenStudy (anonymous):

help plsss

jimthompson5910 (jim_thompson5910):

ignore the pi/3 radian angle for now

jimthompson5910 (jim_thompson5910):

what is the area of the entire circle (radius = 120 ft) ?

OpenStudy (anonymous):

pir squred

jimthompson5910 (jim_thompson5910):

pi*r^2, yes now plug in r = 120 and you get what?

OpenStudy (anonymous):

45238.9

jimthompson5910 (jim_thompson5910):

that's the approximate area, yes

jimthompson5910 (jim_thompson5910):

there are 2pi radians in a full revolution but we only water pi/3 radians in one hour divide the two values: pi/3 divided by 2pi = (pi/3)*(1/(2pi)) = pi/(6pi) = 1/6 this shows us that in one hour, we have watered 1/6th of the entire circle

jimthompson5910 (jim_thompson5910):

so you have to take 1/6 of the area, or divide the area by 6, to get your final answer don't forget to round to the nearest square foot when you're all done

OpenStudy (anonymous):

thank you soo much <3333

jimthompson5910 (jim_thompson5910):

no problem